Problem 4-21 Schedule of cash payments [LO4-2]

The Denver Corporation has forecast the following sales for the first seven months of the year:

January $ 27,000 May $ 27,000
February 29,000 June 33,000
March 31,000 July 35,000
April 37,000

Monthly material purchases are set equal to 25 percent of forecast sales for the next month. Of the total material costs, 30 percent are paid in the month of purchase and 70 percent in the following month. Labor costs will run $5,700 per month, and fixed overhead is $10,500 per month. Interest payments on the debt will be $4,700 for both March and June. Finally, the Denver salesforce will receive a 1.50 percent commission on total sales for the first six months of the year, to be paid on June 30.


Prepare a monthly summary of cash payments for the six-month period from January through June. (Note: Compute prior December purchases to help get total material payments for January.)
Answer #1

Material purchases schedule

December January February March April May June
Purchases 27,000 x 25/100 = 6,750 29,000 x 25/100 = 7,250 31,000 x 25/100 = 7,750 37,000 x 25/100 = 9,250 27,000 x 25/100 = 6,750 33,000 x 25/100 = 8,250 35,000 x 25/100 = 8,750

Schedule of material payment

Month Purchases January February March April May June
December 6,750 6,750 x 70% = 4,725
January 7,250 7,250 x 30% = 2,175 7,250 x 70% = 5,075
February 7,750 7,750 x 30% = 2,325 7,750 x 70% = 5,425
March 9,250 9,250 x 30% = 2,775 9,250 x 70% = 6,475
April 6,750 6,750 x 30% = 2,025 6,750 x 70% = 4,725
May 8,250 8,250 x 30% = 2,475 8,250 x 70% = 5,775
June 8,750 8,750 x 30% = 2,625
Total 6,900 7,400 8,200 8,500 7,200 8,400

Summary of cash payments for the six-month period from January through June

January February March April May June
Payment for material purchases 6,900 7,400 8,200 8,500 7,200 8,400
Payment for labor cost 5,700 5,700 5,700 5,700 5,700 5,700
Payment for fixed overhead 10,500 10,500 10,500 10,500 10,500 10,500
Payment for interest 0 0 4,700 0 0 4,700
Payment for sales commission 0 0 0 0 0 2,760
Total payments 23,100 23,600 29,100 24,700 23,400 32,060

Total sales for the first 6 months = 27,000 + 29,000 + 31,000 + 37,000 + 27,000 + 33,000

= $184,000

Sales commission = 184,000 x 1.5/100

= $2,760

Kindly comment if you need further assistance. Thanks